Subject: Operating System.

Question 3) Assuming a system has 86 available frames and there are two processes P1 with size 20 and P2 with size 100, set up the equations to determine how many frames will be assigned to P1 and P2 using the proportional allocation frame allocation scheme.
Answer #1

3)

Proportional allocation allocates frames to a process depending on the requirements of the process.

Let 'm' be the total number of available frames.

Let the size of process ' pi ' be 'si '

Let ' S ' be the total size of all the processes, i.e, sum of the sizes of the individual processes.

Allocation to process pi = ( si / S ) * m

Allocation of frames to Process P1 = (20 / 120) * 86 = 14

Allocation of frames to Process P2 = (100 / 120) * 86 = 71